# The tame Deligne-Simpson problem

### v1

@@ 4.1. Dimension vectors of $\boldsymbol{\theta}$-stable representations
The Deligne-Simpson problem can be regarded as the trivial stability case of the more general question about the existence of $\boldsymbol{\theta}$-stable representations. The techniques that we have used so far work in this generality as well. Let $Q$ be a star-shaped quiver defined by some tuple of integers $\underline{\nu}$ as in §3.1. Fix $\mathbf{q}\in(\mathbb{C}^{\ast})^{Q_{0}}$ and $\boldsymbol{\theta}\in\mathbb{R}^{Q_{0}}$, define $R_{\mathbf{q},\boldsymbol{\theta}}^{+}:=\{\mathbf{d}\in R^{+}|\mathbf{q}^{\mathbf{d}}=1,\leavevmode\nobreak\ \boldsymbol{\theta}\cdot\mathbf{d}=0\}$ and
@@
$\mathcal{M}_{\boldsymbol{\theta}}(\mathbf{q},\mathbf{d})\stackrel{{\scriptstyle\mathbf{Var}_{0}}}{{\longleftarrow}}\mathcal{M}_{\tilde{\boldsymbol{\theta}}}(\mathbf{q},\mathbf{d})$
@@
(4.1.1) $\mathcal{M}_{B}(\mathbf{d},\boldsymbol{\beta},\boldsymbol{\xi})\stackrel{{\scriptstyle\mathbf{NH}_{1}}}{{\rightarrow}}\mathcal{M}_{Dol}(\mathbf{d}^{\prime},\boldsymbol{\alpha}^{\prime},(\mathcal{O}^{\prime}_{j})_{j})\stackrel{{\scriptstyle\mathbf{Var}_{2}}}{{\leftarrow}}\mathcal{M}_{Dol}(\mathbf{d},\boldsymbol{\alpha},(\mathcal{O}_{j})_{j})\stackrel{{\scriptstyle\mathbf{NH}_{2}}}{{\leftarrow}}\mathcal{M}_{B}(\mathbf{d},\boldsymbol{\beta},\boldsymbol{\xi}^{\prime})\stackrel{{\scriptstyle\mathbf{Iso}_{3}}}{{\rightarrow}}\mathcal{M}_{B}(\mathcal{C}^{\prime}).$
@@ 4.2. Connectedness of character varieties with nongeneric monodromies
As in the proof of Theorem 4.1, we may assume that $\theta_{[j,i]}\geq 0$ for any $j$ and any $i>0$. The vector $\mathbf{d}$ as in the statement of the proposition is called $q$-indivisible in [ST22]. By [ST22, Theorem 6.23], for an almost generic $\tilde{\boldsymbol{\theta}}$ in a small neighbourhood of $\boldsymbol{\theta}$, variation of stability defines a resolution $\mathbf{Var}_{0}:\mathcal{M}_{\tilde{\boldsymbol{\theta}}}(\mathbf{q},\mathbf{d})\longrightarrow\mathcal{M}_{\boldsymbol{\theta}}(\mathbf{q},\mathbf{d})$. The stability condition $\tilde{\boldsymbol{\theta}}$ can be chosen in such a way that $\tilde{\theta}_{[j,i]}>0$ as long as $i>0$. Again, we use the morphisms (4.1.1) to show that $\mathcal{M}_{\tilde{\boldsymbol{\theta}}}(\mathbf{q},\mathbf{d})$ is connected. Since $\mathcal{M}_{\tilde{\boldsymbol{\theta}}}(\mathbf{q},\mathbf{d})$ consists of stable representations, the space $\mathcal{M}_{Dol}(\mathbf{d}^{\prime},\boldsymbol{\alpha}^{\prime},(\mathcal{O}^{\prime}_{j})_{j})$ consists of stable parabolic Higgs bundles. We may choose $\boldsymbol{\alpha}$ to be generic so that $\mathcal{M}_{Dol}(\mathbf{d},\boldsymbol{\alpha},(\mathcal{O}_{j})_{j})$ also consists of stable parabolic Higgs bundles. Indeed, our assumptions on $\mathbf{d}$ and $\mathbf{q}$ imply that $(e,\mathbf{d})$ is indivisible by Lemma 3.6, where $e$ is the degree of the underlying vector bundles in these moduli spaces. By Proposition 3.7, $\mathbf{Var}_{2}$ is bijective. Now, the eigenvalues $\boldsymbol{\xi}^{\prime}$ are generic by Lemma 3.10 and our choice of $\boldsymbol{\alpha}$; therefore, $\mathbf{Iso}_{3}$ is an isomorphism by Proposition 3.11. Finally, the morphisms (4.1.1) combined with Theorem 2.5 show that $\mathcal{M}_{\tilde{\boldsymbol{\theta}}}(\mathbf{q},\mathbf{d})$ is connected. ∎
@@ 4.3. Decomposition of character varieties
Let us show that this morphism between irreducible normal varieties is birational. By [ST22, Proposition 2.15], which is the $\boldsymbol{\theta}$-version of [CBS06, Lemma 7.1], the stratum of $\mathcal{M}_{\boldsymbol{\theta}}(\mathbf{q},m^{\prime}\mathbf{d})$ consisting of mutually nonisomorphic $\mathbf{d}$-dimension $\boldsymbol{\theta}$-stable representations is open (and has dimension $2m^{\prime}$). The open subset of $S^{m^{\prime}}\mathcal{M}_{\boldsymbol{\theta}}(\mathbf{q},\mathbf{d})$ consisting of mutually nonisomorphic $\mathbf{d}$-dimension simple representations is in bijection with this open stratum of $\mathcal{M}_{\boldsymbol{\theta}}(\mathbf{q},m^{\prime}\mathbf{d})$; therefore, they are isomorphic by Zariski main theorem, hence birationality.
@@
Recall that $q_{\infty}=d_{\infty}=1$ and $\mathbf{d}|_{Q}=m\boldsymbol{\delta}$ for some $m\geq 2$, and note that $\mathcal{M}(q_{\infty},d_{\infty})$ is a point. Let us show that $\Psi$ is an isomorphism, which is equivalent to our claim that every $\rho\in\mathcal{M}_{\boldsymbol{\theta}}(\mathbf{q},\mathbf{d})$ decomposes as $\rho_{\infty}\oplus\rho_{1}$. Consider this morphism at the level of representation spaces. Denote by $\mu_{\infty}$ the quasi-Hamiltonian moment map associated to the quiver of type ($\mathbf{Aff}^{\infty}$) as in (1.2) and by $\mu$ the map associated to $Q$. Then, taking the direct sum with $\rho_{\infty}$ defines a closed immersion $\mu^{-1}(\mathbf{q}|_{Q})\hookrightarrow\mu_{\infty}^{-1}(\mathbf{q})$. It follows that $\Psi$ is a closed immersion. Since $\mathbf{d}$ is indivisible, Proposition 4.2 shows that $\mathcal{M}_{\boldsymbol{\theta}}(\mathbf{q},\mathbf{d})$ is connected, and thus irreducible. However, both $\mathcal{M}_{\boldsymbol{\theta}}(\mathbf{q}|_{Q},\mathbf{d}|_{Q})$ and $\mathcal{M}_{\boldsymbol{\theta}}(\mathbf{q},\mathbf{d})$ have dimension $2m$, according to Theorem 4.4 and Lemma 2.8; therefore, $\Psi$ is an isomorphism. ∎

### v2

@@ 4.1. Dimension vectors of $\boldsymbol{\theta}$-stable representations
The Deligne-Simpson problem can be regarded as the trivial stability case of the more general question about the existence of $\boldsymbol{\theta}$-stable representations. The techniques that we have used so far work in this generality as well. Let $Q$ be a star-shaped quiver defined by some tuple of integers $\underline{\nu}$ as in §3.1. Fix $\mathbf{q}\in(\mathbb{C}^{\ast})^{Q_{0}}$ and $\boldsymbol{\theta}\in\mathbb{R}^{Q_{0}}$, define $R_{\mathbf{q},\boldsymbol{\theta}}^{+}:=\{\mathbf{d}\in R^{+}|\mathbf{q}^{\mathbf{d}}=1,\penalty 10000\ \boldsymbol{\theta}\cdot\mathbf{d}=0\}$ and
@@
$\mathcal{M}_{\boldsymbol{\theta}}(\mathbf{q},\mathbf{d})\mathrel{{\mathop{\longleftarrow}\limits^{\mathbf{Var}_{0}}}}\mathcal{M}_{\tilde{\boldsymbol{\theta}}}(\mathbf{q},\mathbf{d})$
@@
(4.1.1) $\mathcal{M}_{B}(\mathbf{d},\boldsymbol{\beta},\boldsymbol{\xi})\mathrel{{\mathop{\rightarrow}\limits^{\mathbf{NH}_{1}}}}\mathcal{M}_{Dol}(\mathbf{d}^{\prime},\boldsymbol{\alpha}^{\prime},(\mathcal{O}^{\prime}_{j})_{j})\mathrel{{\mathop{\leftarrow}\limits^{\mathbf{Var}_{2}}}}\mathcal{M}_{Dol}(\tilde{\mathbf{d}},\boldsymbol{\alpha},(\mathcal{O}_{j})_{j})\mathrel{{\mathop{\leftarrow}\limits^{\mathbf{NH}_{2}}}}\mathcal{M}_{B}(\mathbf{d},\boldsymbol{\beta},\boldsymbol{\xi}^{\prime})\mathrel{{\mathop{\rightarrow}\limits^{\mathbf{Iso}_{3}}}}\mathcal{M}_{B}(\mathcal{C}^{\prime}).$
@@ 4.2. Connectedness of character varieties with nongeneric monodromies
[…]
@@ 4.3. Decomposition of character varieties
Let us show that this morphism between irreducible normal varieties is birational. By [ST22, Proposition 2.15], which is the $\boldsymbol{\theta}$-version of [CBS06, Lemma 7.1], the stratum of $\mathcal{M}_{\boldsymbol{\theta}}(\mathbf{q},m^{\prime}\mathbf{d})$ consisting of mutually nonisomorphic $\mathbf{d}$-dimension $\boldsymbol{\theta}$-stable representations is open (and has dimension $2m^{\prime}$). The open subset of $S^{m^{\prime}}\mathcal{M}_{\boldsymbol{\theta}}(\mathbf{q},\mathbf{d})$ consisting of mutually nonisomorphic $\mathbf{d}$-dimension $\boldsymbol{\theta}$-stable representations is in bijection with this open stratum of $\mathcal{M}_{\boldsymbol{\theta}}(\mathbf{q},m^{\prime}\mathbf{d})$; therefore, they are isomorphic by Zariski main theorem, hence birationality.
@@
Recall that $q_{\infty}=d_{\infty}=1$ and $\mathbf{d}|_{Q}=m\boldsymbol{\delta}$ for some $m\geq 2$, and note that $\mathcal{M}(q_{\infty},d_{\infty})$ is a point. Let us show that $\Psi$ is an isomorphism, which is equivalent to our claim that every $\rho\in\mathcal{M}_{\boldsymbol{\theta}}(\mathbf{q},\mathbf{d})$ decomposes as $\rho_{\infty}\oplus\rho_{1}$. Consider this morphism at the level of representation spaces. Denote by $\mu_{\infty}$ the quasi-Hamiltonian moment map associated to the quiver of type ($\mathbf{Aff}^{\infty}$) as in (1.2) and by $\mu$ the map associated to $Q$. Then, taking the direct sum with $\rho_{\infty}$ defines a closed immersion $\mu^{-1}(\mathbf{q}|_{Q})\hookrightarrow\mu_{\infty}^{-1}(\mathbf{q})$. It follows that $\Psi$ is a closed immersion. Since $\mathbf{d}$ is indivisible, Proposition 4.2 shows that $\mathcal{M}_{\boldsymbol{\theta}}(\mathbf{q},\mathbf{d})$ is connected, and thus irreducible. However, both $\mathcal{M}_{\boldsymbol{\theta}}(\mathbf{q}|_{Q},\mathbf{d}|_{Q})$ and $\mathcal{M}_{\boldsymbol{\theta}}(\mathbf{q},\mathbf{d})$ have dimension $2m$; therefore, $\Psi$ is an isomorphism. ∎